OpenStudy (anonymous):

18.034 divided by 0.04

OpenStudy (anonymous):



hartnn (hartnn):

to eliminate the decimal,you can multiply and divide by 100. so it would be \(\frac{18.034}{0.04}=\frac{18.034*100}{0.04*100}=\frac{1803.4}{4}\) can u solve further?

OpenStudy (anonymous):

\[0.04\sqrt{18.034}\]

OpenStudy (anonymous):

if u have a caulater that helps lol